OpenStudy (anonymous):

Solve for z 6=1/6z+8

OpenStudy (anonymous):

\[6 = \frac{ 1 }{ 6 }z + 8\] ?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

well. we can say: \[6 - 8 = \frac{ 1 }{ 6 } z + 8 - 8\]

OpenStudy (anonymous):

like: a = a so a - 8 = a -8 9 = 9 --> 9 - 8 = 9 - 8 >> 1 = 1

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Wait i dont get it i dont think thats correct.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

well, please give me an example that its not CORRECT!

OpenStudy (igreen):

\(6 = \dfrac{ 1 }{ 6 }z + 8\) \(-\dfrac{ 1 }{ 6 }z = 8 - 6\) \(-\dfrac{ 1 }{ 6 }z = 2\) \(z = \dfrac{1}{3}\)

OpenStudy (igreen):

Sorry that's -1/3

OpenStudy (anonymous):

Yea you were incorrect. Can you see what the question was @PFEH.1999?? Thanks @iGreen

OpenStudy (igreen):

Well @PFEH.1999 was right here: \(6 - 8 = \dfrac{ 1 }{ 6 } z + 8 - 8\) Which will give us: \(-2 = \dfrac{ 1 }{ 6 } z\) Multiply: \(z = -\dfrac{1}{3}\)

OpenStudy (anonymous):

these questions can be solved by different ways :)